OpenStudy (anonymous):

what is the solution to the equation? -2+ 1/2z=8

OpenStudy (anonymous):

\[-2+\frac{1}{2z}=8\] \[\frac{-4z+1}{2z}=8\] \[-4z+1=16z\] \[-4z-16z=-1\] \[-20z=-1\] \[z=\frac{1}{20}\]

OpenStudy (anonymous):

I can't type in fractions. So what does Z=?

OpenStudy (anonymous):

z=1/20 \[z=\frac{1}{20}\]

OpenStudy (anonymous):

no dear on my work I can't use any thing but #'s I go to a online school

OpenStudy (anonymous):

so u want the value of z right !

OpenStudy (anonymous):

z = 0.05

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

z = 1/20 1 ÷ 20 = 0.05 z = 0.05

OpenStudy (anonymous):

It said you were incorrect.

OpenStudy (anonymous):

1/20 = 0.05